# Data-Efficient Extremum-Seeking Control Using Kernel-Based Function Approximation

### v1

@@ 3.3 Guaranteeing a decrease in cost during kernel-based update steps
Given an approximation $m_{k}(\cdot)$ of the input-output map $f(\cdot)$ on the basis of a dataset $\mathcal{D}_{k}$ as in (26), it remains to be assessed whether a kernel-based parameter update step (20) can be performed, by studying the bounds $\underline{b}(\theta)$ and $\overline{b}(\theta,\mu)$. As explained in Section 3.1, these bounds are used to determine whether $-\nabla m_{k}(\theta)$ is a descent direction of $f(\theta)$, and to search for a suitable optimizer gain $\mu\in\mathbb{R}_{>0}$ for which the Armijo condition (21) is satisfied. We perform both these steps without requiring additional experiments to be performed. To obtain these bounds, we draw inspiration from the bounds on function values of an unknown function $f\in\mathcal{H}$ from Scharnhorst et al. (2023), that only require a dataset $\mathcal{D}_{k}$, an upper bound $\bar{\delta}$ on perturbations on the evaluations of $f(\cdot)$ as in Assumption 7, and an upper bound $\Gamma$ on $\lVert f\rVert$ as in Assumption 8. Before presenting the expressions for the bounds $\underline{b}(\theta)$ and $\overline{b}(\theta,\mu)$, for a dataset $\mathcal{D}_{k}$ we define the symmetric matrix
@@
with $K_{\Omega}(\cdot)$ a row vector of kernel functions centered at the inputs $\Omega=\{\widetilde{\theta}_{1},\dots,\widetilde{\theta}_{N_{k}}\}$ as in (29), and where
@@
Given the definitions (34) and (36), the following theorem states that each bound $\underline{b}(\theta)$ and $\overline{b}(\theta,\mu)$ can be evaluated by solving a second-order cone program (SOCP) for a given input $\theta\in\mathbb{R}^{n_{\theta}}$ and optimizer gain $\mu\in\mathbb{R}_{>0}$.
@@
Suppose Assumptions 7 and 8 hold. Let a dataset $\mathcal{D}_{k}=\{(\widetilde{\theta}_{i},\widetilde{y}_{i}):i=1,\dots,N_{k}\}$ of constant inputs $\widetilde{\theta}_{i}$ and corresponding output measurements $\widetilde{y}_{i}$, e.g., as in (14) and (15), respectively, be given. Furthermore, let $m_{k}(\cdot)$ be an approximation of the input-output map $f(\cdot)$ on the basis of this dataset $\mathcal{D}_{k}$ as in (26). Then the minimum value that $\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$ can obtain at a given input $\theta\in\mathbb{R}^{n_{\theta}}$ can be found by solving the SOCP
$\underline{b}(\theta)=$
@@
$\lvert\xi_{i}^{\intercal}\mathfrak{K}(\theta)\beta-\widetilde{y}_{i}\rvert\leq\overline{\delta}\quad\forall i\in\{1,\dots,N_{k}\},$ (38c)
@@
Moreover, for any given $\mu\in\mathbb{R}_{>0}$ and $c\in(0,1)$, the maximum value that $f(\theta^{+})-f(\theta)+c\mu\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$ with $\theta^{+}$ as in (20) can obtain at a given input $\theta\in\mathbb{R}^{n_{\theta}}$ can be found by solving the SOCP
$\overline{b}(\theta,$ $\mu)=$
@@
$\lvert\xi_{i}^{\intercal}\mathfrak{K}^{\prime}(\theta,\theta^{+})\beta-\widetilde{y}_{i}\rvert\leq\overline{\delta}\quad\forall i\in\{1,\dots,N_{k}\}$ (39c)
@@
With the bounds $\underline{b}(\theta)$ and $\overline{b}(\theta,\mu)$ as in Theorem 10, we are ready to state in Algorithm 2 the proposed kernel-based extremum-seeking control algorithm that was briefly outlined in Section 3.1 (see also Figure 1). In this algorithm, we perform the backtracking line search over $\mu\in[\underline{\mu},\overline{\mu}]$ with $\underline{\mu}$ and $\overline{\mu}$ a positive lower and upper bound, respectively, by multiplying $\mu$ by a reduction factor $\varrho\in(0,1)$ until the Armijo condition (21) is guaranteed to be satisfied or $\mu<\underline{\mu}$.
@@
1:Waiting time $T$ and number of experiments $n_{v}$ for a single standard parameter update step as in Algorithm 1; upper bound $\bar{\delta}$ on the measurement perturbations as in Assumption 7; upper bound $\Gamma$ on $\lVert f\rVert$ as in Assumption 8; backtracking line search lower bound $\underline{\mu}$, upper bound $\overline{\mu}$ and reduction factor $\varrho\in(0,1)$; control parameter $c\in(0,1)$ as in the Armijo condition (21); maximum number of update steps $\bar{k}$; initial optimization algorithm state $\widehat{\theta}_{0}$; and a dataset $\mathcal{D}_{0}$ that is either empty, or contains previously obtained input-output data pairs $(\widetilde{\theta},\widetilde{y})$, e.g., as in (14)-(15).
@@
5:     Perform a standard parameter update step as described in Algorithm 1 to update the optimizer state $\widehat{\theta}_{0}$.
6:     Add data collected during the update step to a dataset $\mathcal{D}_{1}$, i.e., $\mathcal{D}_{1}\leftarrow\{(\widetilde{\theta}_{i},\widetilde{y}_{i}):i=1,\dots,n_{v}\}$.
@@
11:     if $-\nabla m_{k}(\widehat{\theta}_{k})$ is guaranteed to be a descent direction of $f(\cdot)$, i.e., if $\underline{b}(\widehat{\theta}_{k})>0$ with $\underline{b}(\cdot)$ as in (38) (cf. (22)) then
12:          Perform a backtracking line search over $\mu\in[\underline{\mu},\overline{\mu}]$, i.e., set $\mu\leftarrow\overline{\mu}$ and
13:          while $\mu\geq\underline{\mu}$ and $\overline{b}(\widehat{\theta}_{k},\mu)>0$ with $\overline{b}(\cdot,\mu)$ as in (39)
@@
16:          if the Armijo condition (21) is guaranteed to be satisfied, i.e., if $\overline{b}(\widehat{\theta}_{k},\mu)\leq 0$ (cf. (23)) then
17:              Perform a kernel-based update (20) without performing additional experiments, i.e., $\widehat{\theta}_{k+1}\leftarrow\widehat{\theta}_{k}-\mu\nabla m_{k}(\widehat{\theta}_{k})$, $\mathcal{D}_{k+1}\leftarrow\mathcal{D}_{k}$, and $N_{k+1}\leftarrow N_{k}$.
@@
22:          Perform a standard parameter update step as described in Algorithm 1 to update the optimizer state $\widehat{\theta}_{k}$.
23:          Add data collected during the update step to a dataset $\mathcal{D}_{k+1}$, i.e., $\mathcal{D}_{k+1}\leftarrow\mathcal{D}_{k}\cup\{(\widetilde{\theta}_{i},\widetilde{y}_{i}):i=N_{k}+1,\dots,N_{k}+n_{v}\}$ and $N_{k+1}\leftarrow N_{k}+n_{v}$.
@@ 4 Stability analysis for kernel-based extremum-seeking control
In this section, we perform a stability analysis for the kernel-based extremum-seeking approach described in Algorithm 2. To this end, we define for any input $\theta\in\mathbb{R}^{n_{\theta}}$ and any minimizer $\theta^{*}\in\mathcal{C}$ of $f(\theta)$ the function
$V(\theta):=f(\theta)-f(\theta^{*}).$ (40)
Furthermore, given this $V(\cdot)$, we define the function (see also Hazeleger et al. (2022, Equation (22)))
@@
with $\eta\in\mathbb{R}_{>0}$ a constant, $x$ the system state, and $\psi$ a memory state for the optimizer state at the start of the last standard update step. That is, $\widetilde{\theta}_{N_{k}}=\psi+v_{n_{v}}(\psi)$ is the last input that has been applied to the system (cf. (14) and Lines 5 and 22 of Algorithm 2). Note that in (41), $V_{p}(\psi,x)$ contains terms relating to the distance of the system state to the attractor $\mathcal{A}(\psi+v_{n_{v}}(\psi))$ (cf. Assumption 2(i)) corresponding to the last applied input $\psi+v_{n_{v}}(\psi)$, and the distance from the corresponding optimizer state $\psi$ to the set of minimizers $\mathcal{C}$ (cf. (42)). That is, $V_{p}(\psi,x)$ is small if the system state $x$ is close to the attractor corresponding to the last applied input and the corresponding optimizer state $\psi$ is close to the set of minimizers $\mathcal{C}$, whereas $V(\theta)$ in (41) relates to the distance of the current optimizer state $\theta$ to the set of minimizers $\mathcal{C}$ since $\omega_{1}(\lVert\theta\rVert_{\mathcal{C}})\leq V(\theta)\leq\omega_{2}(\lVert\theta\rVert_{\mathcal{C}})$ (cf. (40) and Assumption 10). Thus, $W(\psi,x,\theta)$ is small when both the current optimizer state $\theta$ and memory state $\psi$ are close to $\mathcal{C}$, and $x$ is close to the attractor corresponding to the last applied input $\psi+v_{n_{v}}(\psi)$.
@@
Suppose that the system $\Sigma_{p}$ in Definition 1 satisfies Assumption 2, and that the optimization algorithm $\Sigma$ in (8) satisfies Assumption 4. Let $x$ and $\theta$ denote, respectively, the state of $\Sigma_{p}$ and $\Sigma$ at the start of the current algorithm update step $k$. Furthermore, let $\psi$ denote a memory state for optimizer state at the start of the last standard update step, i.e., $\widetilde{\theta}_{N_{k}}=\psi+v_{n_{v}}(\psi)$ is the last input that has been applied to the system (cf. (14)). Then, for any $\Delta_{\theta},\Delta_{x}\in\mathbb{R}_{>0}$ such that $\lVert\theta\rVert_{\mathcal{C}}\leq\Delta_{\theta}$, $\lVert\psi\rVert\leq\Delta_{\theta}$, and $\lVert x\rVert_{\mathcal{A}(\psi+v_{n_{v}}(\psi))}\leq\Delta_{x}$, there exists a sufficiently long waiting time $T^{*}\geq T$, with $T$ as in Assumption 2(vi), and a class-$\mathcal{K}_{\infty}$ function $\widetilde{\gamma}(\cdot)$, such that for a standard parameter update step as described in Algorithm 1 the upper bound
@@
Suppose Assumptions 2(iii), 10, 7 and 8 hold. Let $m_{k}(\cdot)$ be an approximation of $f(\cdot)$ obtained as in (26) from a dataset $\mathcal{D}_{k}=\{\widetilde{\theta}_{i},\widetilde{y}_{i}:i=1,\dots,N_{k}\}$ with constant inputs $\widetilde{\theta}_{i}$ and their corresponding output measurements $\widetilde{y}_{i}$, e.g., as in (14) and (15), respectively. Furthermore, let $\theta^{+}$ be the result of a kernel-based update step as in (20). Finally, let $\underline{b}(\theta)$ and $\overline{b}(\theta,\mu)$ be bounds as in Theorem 10. If, for any $\theta\in\mathbb{R}^{n_{\theta}}\backslash\mathcal{C}$, $\underline{b}(\theta)>0$ and $\mu\in\mathbb{R}_{>0}$ in (20) is chosen such that $\overline{b}(\theta,\mu)\leq 0$, then for any $\Delta_{\theta}\in\mathbb{R}_{>0}$ such that $\lVert\theta\rVert_{\mathcal{C}}\leq\Delta_{\theta}$, there exists a class-$\mathcal{K}$ function $\widetilde{\rho}_{k}(\cdot)$ such that
@@
Note with respect to Remark 5 and Lemmas 12 and 13, that when $V(\cdot)$ would be chosen as general locally Lipschitz function, instead of as in (40), it is not immediately clear under which conditions a kernel-based update step is ‘helpful’, i.e., under which conditions a kernel-based update step leads to a decrease in $V(\theta)$. By defining $V(\cdot)$ explicitly as in (40), such decrease can be guaranteed by satisfaction of the Armijo condition (21) if $-\nabla m_{k}(\theta)$ is a descent direction as shown in the proof of Lemma 13. However, by defining $V(\cdot)$ as in (40), (10) and (11) must hold to satisfy Hazeleger et al. (2022, Assumption 7) used in the proof of Lemma 12, instead of requiring only a locally Lipschitz function $V(\cdot)$ that satisfies both $\omega_{1}(\lVert\theta\rVert_{\mathcal{C}})\leq V(\theta)\leq\omega_{2}(\lVert\theta\rVert_{\mathcal{C}})$ and $\max_{\theta^{+}\in F(\theta,Y(\theta))}V(\theta^{+})-V(\theta)\leq-\rho(V(\theta))+\delta$.
@@
Suppose that the system $\Sigma_{p}$ in Definition 1 and optimizer $\Sigma$ in (8) satisfy Assumptions 2, 4, 7, and 8. For any $\theta_{0},\psi_{0}\in\mathbb{R}^{n_{\theta}}$ and $x_{0}\in\mathcal{X}$ with $\lVert\theta_{0}\rVert_{\mathcal{C}}\leq\Delta_{\theta}$, $\lVert\psi_{0}\rVert_{\mathcal{C}}\leq\Delta_{\theta}$ and $\lVert x_{0}\rVert_{\mathcal{A}(\psi_{0}+v_{n_{v}}(\psi_{0}))}\leq\Delta_{x}$ for some $\Delta_{\theta},\Delta_{x}\in\mathbb{R}_{>0}$, there exists a sufficiently long waiting time $T^{*}\geq T$ for standard parameter updates as in Algorithm 1, with $T$ satisfying Assumption 2(vi), and a class-$\mathcal{K}_{\infty}$ function $\varphi(\cdot)$, such that for the kernel-based extremum-seeking algorithm described in Algorithm 2 the optimizer state $\theta$ converges to the set
$\left\{\theta\in\mathbb{R}^{n_{\theta}}:\lVert\theta\rVert_{\mathcal{C}}\leq\omega_{1}^{-1}\left(\frac{1}{2}\varphi(\gamma+2\delta+\delta_{V})\right)\right\}$ (47)
@@
Note that the input applied to the system $\Sigma_{p}$ is not changed during kernel-based update steps (cf. Algorithm 2, Line 17), and that kernel-based update steps are instantaneous (since there is no waiting time $T$ for kernel-based update steps). As a consequence, the memory state $\psi$ and system state $x$ do not change during kernel-based update steps. Therefore, while the optimizer state $\theta$ converges to a neighborhood of the set of minimizers as stated in Theorem 15, the system output $y$ does not necessarily converge of a neighborhood of the minimum of the input-output map $f(\cdot)$. However, by Assumptions 2(i) and 2(ii), this discrepancy between the value of the steady-state input-output map associated with the current optimizer state and the real plant output can always be made arbitrarily small by applying the optimizer state as an input to the plant and waiting sufficiently long (after the last kernel-based update step).

### v2

@@ 3.3 Guaranteeing a decrease in cost during kernel-based update steps
Given an approximation $m_{k}(\cdot)$ of the input-output map $f(\cdot)$ on the basis of a dataset $\mathcal{D}_{k}$ as in (26), it remains to be assessed whether a kernel-based parameter update step (19) can be performed. To this end, as explained in Section 3.1, we need to determine whether $-\nabla m_{k}(\theta)$ is a descent direction of $f(\theta)$, and to search for a suitable optimizer gain $\mu\in\mathbb{R}_{>0}$ for which the Armijo condition (20) is satisfied. We perform both these steps without requiring additional measurements to be performed, by studying the bounds $\underaccent{\bar}{b}(\theta)$ and $\bar{b}(\theta,\mu)$ in (22) and (23). To construct expressions for these bounds, we define the symmetric matrix
@@
with $K_{\Omega}(\cdot)$ a row vector of kernel functions centered at the inputs $\Omega=\{\widetilde{\theta}_{1},\widetilde{\theta}_{2},\dots,\widetilde{\theta}_{N_{k}}\}$ as in (27), and where
@@
Given the definitions (32) and (34), the following theorem states that each bound $\underaccent{\bar}{b}(\theta)$ and $\bar{b}(\theta,\mu)$ can be evaluated by solving a second-order cone program (SOCP) for a given input $\theta\in\mathbb{R}^{n_{\theta}}$ and optimizer gain $\mu\in\mathbb{R}_{>0}$.
@@
Suppose Assumptions 7 and 8 hold. Let a dataset $\mathcal{D}_{k}=\{(\widetilde{\theta}_{i},\widetilde{y}_{i})\mid i=1,2,\dots,N_{k}\}$ of constant inputs $\widetilde{\theta}_{i}$ and corresponding output measurements $\widetilde{y}_{i}$, e.g., as in (10) and (11), respectively, be given. Furthermore, let $m_{k}(\cdot)$ be an approximation of the input-output map $f(\cdot)$ on the basis of this dataset $\mathcal{D}_{k}$ as in (26). Then the minimum value that $\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$ can attain at a given input $\theta\in\mathbb{R}^{n_{\theta}}$ can be found by solving the SOCP
$\underaccent{\bar}{b}(\theta)=$
@@
$\lvert\xi_{i}^{\intercal}\mathfrak{K}(\theta)\beta-\widetilde{y}_{i}\rvert\leq\bar{\delta}\quad\forall i\in\{1,2,\dots,N_{k}\},$ (36c)
@@
Moreover, for any given $\mu\in\mathbb{R}_{>0}$ and $c\in(0,1)$, the maximum value that $f(\theta^{+})-f(\theta)+c\mu\nabla f(\theta)^{\intercal}\nabla m_{k}(\theta)$ with $\theta^{+}$ as in (19) can attain at a given input $\theta\in\mathbb{R}^{n_{\theta}}$ can be found by solving the SOCP
$\bar{b}(\theta,$ $\mu)=$
@@
$\lvert\xi_{i}^{\intercal}\mathfrak{K}^{\prime}(\theta,\theta^{+})\beta-\widetilde{y}_{i}\rvert\leq\bar{\delta}\;\forall i\in\{1,2,\dots,N_{k}\}$ (37c)
@@
With the bounds $\underaccent{\bar}{b}(\theta)$ and $\bar{b}(\theta,\mu)$ as in Theorem 10, we state in Algorithm 1 the proposed kernel-based extremum-seeking control algorithm that we outlined briefly in Section 3.1 (see also Figure 1). In this algorithm, we perform the backtracking line search over $\mu\in[\underaccent{\bar}{\mu},\bar{\mu}]$ with $\underaccent{\bar}{\mu}$ and $\bar{\mu}$ a positive lower and upper bound, respectively, by multiplying $\mu$ by a reduction factor $\varrho\in(0,1)$ until the Armijo condition (20) is guaranteed to be satisfied or $\mu<\underaccent{\bar}{\mu}$.
@@
1:Waiting time $T$ and number of measurements $n_{v}$ for a single standard parameter update step as in (10)-(13); upper bound $\bar{\delta}$ on the measurement perturbations as in Assumption 7; upper bound $\Gamma$ on $\lVert f\rVert$ as in Assumption 8; backtracking line search lower bound $\underaccent{\bar}{\mu}$, upper bound $\bar{\mu}$ and reduction factor $\varrho\in(0,1)$; control parameter $c\in(0,1)$ as in the Armijo condition (20); maximum number of update steps $\bar{k}$; initial optimization algorithm state $\widehat{\theta}_{0}$; and a dataset $\mathcal{D}_{0}$ that is either empty, or contains previously obtained input-output data pairs $(\widetilde{\theta},\widetilde{y})$, e.g., as in (10)-(11).
@@
5:     Perform a standard parameter update step as in (10)-(13) to update the optimizer state $\widehat{\theta}_{0}$.
6:     Add data collected during the update step to a dataset $\mathcal{D}_{1}$, i.e., $\mathcal{D}_{1}\leftarrow\{(\widetilde{\theta}_{i},\widetilde{y}_{i})\mid i=1,2,\dots,n_{v}\}$.
@@
11:     if $-\nabla m_{k}(\widehat{\theta}_{k})$ is guaranteed to be a descent direction of $f(\cdot)$, i.e., if $\underaccent{\bar}{b}(\widehat{\theta}_{k})>0$ with $\underaccent{\bar}{b}(\cdot)$ as in (36) (cf. (22)) then
12:          Perform a backtracking line search over $\mu\in[\underaccent{\bar}{\mu},\bar{\mu}]$, i.e., set $\mu\leftarrow\bar{\mu}$ and
13:          while $\mu\geq\underaccent{\bar}{\mu}$ and $\bar{b}(\widehat{\theta}_{k},\mu)>0$ with $\bar{b}(\cdot,\mu)$ as in (37)
@@
16:          if the Armijo condition (20) is guaranteed to be satisfied, i.e., if $\bar{b}(\widehat{\theta}_{k},\mu)\leq 0$ (cf. (23)) then
17:              Perform a kernel-based update (19) without performing additional measurements, i.e., $\widehat{\theta}_{k+1}\leftarrow\widehat{\theta}_{k}-\mu\nabla m_{k}(\widehat{\theta}_{k})$, $\mathcal{D}_{k+1}\leftarrow\mathcal{D}_{k}$, and $N_{k+1}\leftarrow N_{k}$.
@@
22:          Perform a standard parameter update step as in (10)-(13) to update the optimizer state $\widehat{\theta}_{k}$.
23:          Add data collected during the update step to a dataset $\mathcal{D}_{k+1}$, i.e., $\mathcal{D}_{k+1}\leftarrow\mathcal{D}_{k}\cup\{(\widetilde{\theta}_{i},\widetilde{y}_{i})\mid i=N_{k}+1,\dots,N_{k}+n_{v}\}$ and $N_{k+1}\leftarrow N_{k}+n_{v}$.
@@ 4 Stability analysis for kernel-based extremum-seeking control
In this section, we perform a stability analysis for the kernel-based extremum-seeking approach described in Algorithm 1. To this end, we define the function (see also Hazeleger et al. (2022, Equation (22)))
@@
with $\eta\in\mathbb{R}_{>0}$ a constant, $x$ the system state, and $\psi$ a memory state for the optimizer state at the start of the last standard update step. That is, $\widetilde{\theta}_{N_{k}}=\psi+v_{n_{v}}(\psi)$ is the last input that has been applied to the system (cf. (10) and Lines 5 and 22 of Algorithm 1). Note that in (38), $V_{p}(\psi,x)$ contains terms relating to the distance of the system state to the attractor $\mathcal{A}(\psi+v_{n_{v}}(\psi))$ (cf. Assumption 2(i)) corresponding to the last applied input $\psi+v_{n_{v}}(\psi)$, and the distance from the corresponding optimizer state $\psi$ to the set of minimizers $\mathcal{C}$ (cf. (39)). That is, $V_{p}(\psi,x)$ is small if the system state $x$ is close to the attractor corresponding to the last applied input and the corresponding optimizer state $\psi$ is close to the set of minimizers $\mathcal{C}$, whereas $V(\theta)$ in (38) relates to the distance of the current optimizer state $\theta$ to the set of minimizers $\mathcal{C}$ since $\omega_{1}(\lVert\theta\rVert_{\mathcal{C}})\leq V(\theta)\leq\omega_{2}(\lVert\theta\rVert_{\mathcal{C}})$ (cf. Assumption 14). Thus, $W(\psi,x,\theta)$ is small when both the current optimizer state $\theta$ and memory state $\psi$ are close to $\mathcal{C}$, and $x$ is close to the attractor corresponding to the last applied input $\psi+v_{n_{v}}(\psi)$.
@@
Suppose that the system $\Sigma_{p}$ in Definition 1 satisfies Assumption 2, and that the optimization algorithm $\Sigma$ in (8) satisfies Assumption 4. Let $x$ and $\theta$ denote, respectively, the state of $\Sigma_{p}$ and $\Sigma$ at the start of the current algorithm update step $k$. Furthermore, let $\psi$ denote a memory state for optimizer state at the start of the last standard update step, i.e., $\widetilde{\theta}_{N_{k}}=\psi+v_{n_{v}}(\psi)$ is the last input that has been applied to the system (cf. (10)). Then, for any $\Delta_{\theta},\Delta_{x}\in\mathbb{R}_{>0}$ such that $\lVert\theta\rVert_{\mathcal{C}}\leq\Delta_{\theta}$, $\lVert\psi\rVert\leq\Delta_{\theta}$, and $\lVert x\rVert_{\mathcal{A}(\psi+v_{n_{v}}(\psi))}\leq\Delta_{x}$, there exists a sufficiently long waiting time $T^{*}\geq T$, with $T$ as in Assumption 2(vi), and a class-$\mathcal{K}_{\infty}$ function $\widetilde{\gamma}(\cdot)$, such that for a standard parameter update step as in (10)-(13) the upper bound
@@
Suppose Assumptions 2(iii), 14, 7 and 8 hold. Let $m_{k}(\cdot)$ be an approximation of $f(\cdot)$ obtained as in (26) from a dataset $\mathcal{D}_{k}=\{\widetilde{\theta}_{i},\widetilde{y}_{i}\mid i=1,2,\dots,N_{k}\}$ with constant inputs $\widetilde{\theta}_{i}$ and their corresponding output measurements $\widetilde{y}_{i}$, e.g., as in (10) and (11), respectively. Furthermore, let $\theta^{+}$ be the result of a kernel-based update step as in (19). Finally, let $\underaccent{\bar}{b}(\theta)$ and $\bar{b}(\theta,\mu)$ be bounds as in Theorem 10. If, for any $\theta\in\mathbb{R}^{n_{\theta}}\backslash\mathcal{C}$, $\underaccent{\bar}{b}(\theta)>0$ and $\mu\in\mathbb{R}_{>0}$ in (19) is chosen such that $\bar{b}(\theta,\mu)\leq 0$, then for any $\Delta_{\theta}\in\mathbb{R}_{>0}$ such that $\lVert\theta\rVert_{\mathcal{C}}\leq\Delta_{\theta}$, there exists a class-$\mathcal{K}$ function $\widetilde{\rho}_{k}(\cdot)$ such that
@@
Note with respect to Remark 5 and Lemmas 12 and 13, that when $V(\cdot)$ would be chosen as general locally Lipschitz function, instead of as in (16), it is not immediately clear under which conditions a kernel-based update step is ‘helpful’, i.e., under which conditions a kernel-based update step leads to a decrease in $V(\theta)$. By defining $V(\cdot)$ explicitly as in (16), such decrease can be guaranteed by satisfaction of the Armijo condition (20) if $-\nabla m_{k}(\theta)$ is a descent direction as shown in the proof of Lemma 13.
@@
Suppose that the system $\Sigma_{p}$ in Definition 1 and optimizer $\Sigma$ in (8) satisfy Assumptions 2, 4, 7, and 8. For any $\theta_{0},\psi_{0}\in\mathbb{R}^{n_{\theta}}$ and $x_{0}\in\mathcal{X}$ with $\lVert\theta_{0}\rVert_{\mathcal{C}}\leq\Delta_{\theta}$, $\lVert\psi_{0}\rVert_{\mathcal{C}}\leq\Delta_{\theta}$ and $\lVert x_{0}\rVert_{\mathcal{A}(\psi_{0}+v_{n_{v}}(\psi_{0}))}\leq\Delta_{x}$ for some $\Delta_{\theta},\Delta_{x}\in\mathbb{R}_{>0}$, there exists a sufficiently long waiting time $T^{*}\geq T$ for standard parameter updates as in (10)-(13), with $T$ satisfying Assumption 2(vi), and a class-$\mathcal{K}_{\infty}$ function $\varphi(\cdot)$, such that for the kernel-based extremum-seeking algorithm described in Algorithm 1 the optimizer state $\theta$ converges to the set
$\left\{\theta\in\mathbb{R}^{n_{\theta}}\mid\lVert\theta\rVert_{\mathcal{C}}\leq\omega_{1}^{-1}\left(\frac{1}{2}\varphi(\gamma+2\delta+\delta_{V})\right)\right\}$ (44)
@@
Note that the input applied to the system $\Sigma_{p}$ is not changed during kernel-based update steps (cf. Algorithm 1, Line 17), and that the duration of kernel-based update steps is negligible compared to the timescale at which the system dynamics evolve (since there is no waiting time $T$ for kernel-based update steps) and can therefore be considered instantaneous. As a consequence, the memory state $\psi$ and system state $x$ do not change (significantly) during kernel-based update steps. Therefore, while the optimizer state $\theta$ converges to a neighborhood of the set of minimizers as stated in Theorem 15, the system output $y$ does not necessarily converge of a neighborhood of the minimum of the input-output map $f(\cdot)$. However, by Assumptions 2(i) and 2(ii), this discrepancy between the value of the steady-state input-output map associated with the current optimizer state and the real plant output can always be made arbitrarily small by applying the optimizer state as an input to the plant and waiting sufficiently long (after the last kernel-based update step).